OpenStudy (anonymous):

Larry painted forty feet of fencing in two point five hours. At that rate, how many feet of fencing can Larry paint in 8 hours? sixteen feet one hundred twenty feet one hundred twenty-eight feet one hundred sixty feet

OpenStudy (anonymous):

set it as a fractional ratio, simplify and multiply to eliminate uneccessary units \[\frac {40 [ft]}{2.5 [hrs]}=\frac{16[ft]}{1[hr]}\] \[\frac{16[ft]}{1\cancel{[hr]}}\times 8 \cancel{[hrs]}=128ft\]